Stocks A and B have the following historical returns:

Year
2012 -20.10% -12.60%
2013 20.50 24.30
2014 17.25 30.90
2015 -3.25 -10.10
2016 29.50 11.40
  1. Calculate the average rate of return for each stock during the 5-year period. Round your answers to two decimal places.
    Stock A %
    Stock B %
  2. Assume that someone held a portfolio consisting of 50% of Stock A and 50% of Stock B. What would have been the realized rate of return on the portfolio in each year? What would have been the average return on the portfolio during this period? Round your answers to two decimal places.
    Year Portfolio
    2012 %
    2013 %
    2014 %
    2015 %
    2016 %
    Average return %
  3. Calculate the standard deviation of returns for each stock and for the portfolio. Round your answers to two decimal places.
    rA rB Portfolio
    Std. Dev. % % %
Answer #1

A Average rate of return for each stock during 5 years will be aa follows :-

Stock A = -20.10+20.50+17.25+(-3.25)+29.50/5

= 43.90/5 ,= 8.78%

Stock B = -12.60+24.30+30.90+(-10.10)+11.40/5

= 43.9/5 = 8.78%

Stock A 8.78%
Stock B 8.78%

B. Following will be the Return of portfolio each year and average return :-

Year Portfolio
2012 -16.35% (-20.10x0.5+-12.60x0.5)
2013 22.40% (20.50x0.50+24.30x0.50)
2014 24.075% (17.25x0.5+30.90x0.5)
2015 -6.675% (-3.25x0.5+-10.10x0.5)
2016 20.45% (29.50x0.5+11.40x0.5)
Average return 8.78% {(-16.35+22.40+24.075+-6.675+20.45)/5}

C . Standard deviation of returns of each stock and portfolio is as follows

rA rB portfolio
Standard deviation 40.21% 39.38% 37.76%

Working note :-

1. Standard deviation for sotck A

m = average return

Year return (x) x-m (here m = 8.78) (x-m)2
2012 -20.10 -28.88 (-20.10-8.78) 834.05
2013 20.50 11.72 (20.50-8.78) 137.36
2014 17.25 8.47 (17.25-8.78) 71.74
2015 -3.25 -12.03 (-3.25-8.78) 144.72
2016 29.50 20.72 (29.50-8.78) 429.32
Total m = 8.78 1617.19

Standard deviation of stock A = √1617.19 = 40.21%.

Calculation of standard deviation of stock B

Year return (x) (x-m) (x-m)2
2012 -12.60 -21.38 (-12.60-8.78) 457.10
2013 24.30 15.52 (24.30-8.78) 240.87
2014 30.90

22.12 (30.90-8.78)

489.29
2015 -10.10 -18.88 (-10.10-8.78) 356.45
2016 11.40 2.62 (11.40-8.78) 6.86
Total m = 8.78 1550.59

Standard deviation of stock B = √1550.59 = 39.38%.

Calculation of is standard deviation of portfolio

Year return (x) (x-m) (x-m)2
2012 -16.35 -25.13 (-16.35-8.78) 631.52
2013 22.40 13.62 (22.40-8.78) 185.50
2014 24.075 15.295 (24.075-8.78) 233.94
2015 -6.675 -15.454 (-6.675-8.78) 238.83
2016 20.45 11.67 (20.45-8.78) 136.19
total m = 8.78 1425.97

Standard deviation of portfolio = √1425.97 = 37.76%